Session: Jun 2006 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Medicina (Morsiani at Sun Jun 11 15:40:19 2006) Success :

Westerbork (Tony Foley at Mon Jun 12 06:51:32 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems

Effelsberg (A. Polatidis at Mon Jun 12 08:46:30 2006) Success :

Torun (K. Borkowski at Mon Jun 12 08:58:02 2006) Minor failures : RECE
No know problems till 12:09 UT when our receivers again (see N06M2) suffered from power supply failure (at the time of this writing still not repaired, so CL experiments are being postponed).

Onsala (Michael Lindqvist at Mon Jun 12 09:46:43 2006) Minor failures :
Still problem with VC03 (see EB032A). However, here VC03 and VC04 are "copies" of VC01 and VC02 if I understand the schedule.

Noto (P. Cassaro at Sat Jun 17 12:11:34 2006) Success :

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:20:00 2006) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Jun 20 09:20:30 2006) Success :
The Darnhall telescope was used. No known problems.

EVN experiment feedback facility: entire session

Entire Session

Session: Jun 2006 (session 2)
Wavelengths: 30cm 18/21 5 6 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Jun 5 12:32:29 2006) Minor failures :
This note concerns the UHF experiments only. Shortly before the first experiment we could manage to install the Lange coupler to produce two circular polarizations instead of the linear polarization announced in the block schedule. Also, we failed to install the phase cal.

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:14:12 2006) Success : PCAL
The MkII telescope was used (Jb2). During this session the Mk2 phase cal. signal was found to be too high and could not be easily reduced in strength. I therefore switched off the phas cal. signal for the entire session (for Mk2 only).